OpenStudy (anonymous):

Find the exact solution for 3(inversetan)x=pi

OpenStudy (mimi_x3):

\[3arctanx = \pi => arctanx = \frac{\pi}{3} => x= \tan\frac{\pi}{3} = \sqrt{3} \] Do you understand?

OpenStudy (anonymous):

Okay I see, you just try to get tan by itself then use the unit cirlce for the sqrt of 3. thank you!

OpenStudy (mimi_x3):

np